OpenStudy (anonymous):

what is the square root is between 5 and 6

OpenStudy (anonymous):

what are you answer choices?

OpenStudy (anonymous):

I'll give you a hint. \[5^2<x^2<6^2\]

OpenStudy (anonymous):

root 29,21,40,48

OpenStudy (anonymous):

which one is it in that choice

OpenStudy (anonymous):

Look at my hint.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Hurry!!!

OpenStudy (anonymous):

\[\sqrt{5^2}<\sqrt{x^2}<\sqrt{6^2}\] what is 5^2 and 6^2?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

...

OpenStudy (anonymous):

5 x 5 and 6 x6 what are they, respectively?

OpenStudy (anonymous):

i got it 61

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

what is 5 x 5?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

yes. And 6 x 6?

OpenStudy (anonymous):

36

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Now we know that \[\sqrt{25}<\sqrt x<\sqrt{36}\]

OpenStudy (anonymous):

which of your choices is between 25-36?

OpenStudy (anonymous):

29

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

ok thanks @monokerous

OpenStudy (anonymous):

thanks for helping

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

Which of the following square roots is between 6 and 7?